### v1

# Triplet of kindred prompt-2 emitters in mass-8 proton-rich nuclei

Abstract
A triplet of kindred prompt-2 emitters in =8 nuclei has been demonstrated. Two of these are the ground state of 8C and its isobaric analog state in 8B, both of which are analogs of the halo or thick-skinned nucleus 8He. The third member is the recently found fourth 1+ state in 8B. This new 8B state at =8.4 MeV was observed to decay to the ground state of 6Li by 2 emission. Momentum correlations between the decay products indicate that it is not a sequential 2 decay through a 7Be intermediate state, but indicative of prompt 2 emission with correlations similar to those of the other members of the triplet. Ab initio calculations with the symmetry-adapted no-core shell model indicate that these three states have very similar spatial wavefunctions, but with nucleons coupling to different spins, isospins, and isospin projections. The triplet of 2 emitters all have oblate shapes and by stripping two protons equatorially, decay to states which have prolate shapes.
As one goes away from the valley of stability on the nuclear chart one finds exotic nuclei that have neutron halos, on the neutron-rich side, and decay by the simultaneous emission of two protons, on the proton-rich side. These halo systems and the prompt 2 emitters can sometimes be associated with similar nuclear structure. For instance, the mirror of the two-neutron-halo nucleus 6He is 6Be which is not a two-proton halo system as the two valence protons are unbound, but these two protons are emitted simultaneously [Egorova:2012]. Another example is 8He, often described as having a 4 halo [AlKhalili:2004], and its mirror 8C decays by two sequential steps of 2 emission [Charity:2010]. These two nuclei are the end members of an isospin multiplet which connects a series of 0+, =2 states with =8 (see Fig. 1) with similar exotic structure.
Prompt 2 decay is an exotic decay mode that is often found for ground states of even- isotopes just beyond the proton drip line [Pfutzner:2023]. For the heavier ground-state 2 emitters, this decay represents the only energy-conserving particle-decay mode. For lighter 2 emitters, there is often an energetically-allowed 1 decay process through an intermediate state which is very broad (short lived) blurring the distinction between prompt and sequential 2 decay.
The isobaric analog states of the prompt-2-decaying ground states of 8C and 12O, respectively in 8B (Fig. 1) and 12N, also undergo prompt 2 decay [Charity:2010, Brown:2014, Webb:2019a]. However, these decays are also forced by conservation laws as they represent the only particle decay mode that conserves both energy and isospin.
Examples of 2 decay from excited states of isotopes inside the drip line where such restrictions from conservation laws are typically absent, are rare and must result from some peculiar nuclear structure favoring 2 emission over 1 decay. One such example is the second 0+ state in 10C [Charity:2022] which is the mirror of a highly-deformed -cluster configuration in 10Be [Freer:2006]. In this case there is at present no theoretical understanding for the strength of the 2 decay.
In the present work, we find another case of 2 decay with competing 1 decay channels, specifically the 1 state in 8B, however in this case ab initio nuclear-structure calculations provide an explanation for the exotic decay. This 8B state should be part of a =1+, =1 isospin multiplet located in energy just below the =2 multiplet (Fig. 1). These =1 states have similar exotic structure to the =2 states to which they are related largely by a spin-flip of one of the nucleons. The proton-rich members (red levels) of the =0+ = 2 and =1+ =1 multiplets comprise a triplet of 2 emitters with similar decay correlations.
Experimental Data.— The 0 ground state of 8C was first observed in missing-mass spectroscopy [Robertson:1974, Tribble:1976, Robertson:1976] and, from its measured mass, is found to be unstable for decay into the 4+ exit channel. This exit channel was first investigated with invariant-mass spectroscopy where 8C resonances were created via neutron knockout from an =69 MeV 9C beam by a 9Be target [Charity:2010, Charity:2011]. The decay-energy spectrum from that work is shown in Fig. 2(a) where the narrow 0 ground-state peak sits upon little background. Correlations between the momentum vectors of the decay fragments indicate that this state decays by two steps of prompt 2 emission [Charity:2010, Charity:2011]. It is the first step, the prompt 2 decay of 8C to the 0 ground state of 6Be, that is of importance for the present work.
With the same reaction, but knocking out a proton instead of a neutron, 2 emission was also found to occur in 8B using invariant-mass spectroscopy of the 2+6Li exit channel [Charity:2010, Brown:2014]. Figure 2(b) shows the 8B excitation energy calculated assuming that the decay populates 6Li in its 1+ ground state. This assumption is incorrect for the narrow low-energy peak as the coincidence Doppler-corrected -ray energy spectrum, shown by the inset, shows the clear signature of the 3.562-MeV ray (full absorption and first-escape peaks indicated by the two arrows) depopulating the 0 isobaric analog state in 6Li [Brown:2014]. Thus the true excitation energy of this 2-emitting state should be increased by the -ray energy. This makes the excitation energy of this state 10.619 MeV, the known excitation energy of 0, the isobaric analog of the ground state of 8C [ENSDF]. The decays of the =2 analog states [8C(0) and 8B()] are illustrated in Fig. 3. Both of these states 2 decay to =1 analog states [6Be(0) and 6Li(0)].
In this work we consider the nature of the second peak in Fig. 2 at =8.4 MeV. This peak is associated with a state of lower excitation energy than that for the narrow 0 peak as it decays to the ground state of 6Li [CharityLong:2025], i.e. there is no coincidence ray and thus =. The existence of this state was confirmed from a second experiment, where the fragmentation products from an =65-MeV 13O beam with a 9Be target [CharityLong:2025] are analyzed with the invariant-mass method. The weighted mean resonance parameters, from the two data sets, are =8.40(4) MeV and =0.88(7) MeV [CharityLong:2025].
The level scheme for the decay of the 8.4-MeV state, shown in Fig. 3(b), indicates the final products can be produced through either one step of prompt 2 decay step or two sequential steps of 1 decay.
For the latter, there is only one known proton-decaying state in 7Be (5/2) through which the sequential decay could proceed.
Information on the nature of the 2 decay can be gleaned from the momentum correlations between decay fragments.
Figures 4(b) and 4(c) show the decay correlations for the three states of interest in terms of the fraction of the decay energy in the - and -core sub-systems. (The core here refers to the daughter, i.e., 6Li[1+, T = 0], 6Li[0+, T = 1], and 6Be[0+, T = 1], from bottom to top in Fig. 3.) The results for 8C(0) and 8B(0) (red and green data points) are from [Charity:2011, Brown:2014]. Background-subtracted correlations for the 8.4-MeV state were extracted in Ref. [CharityLong:2025] from the two invariant-mass data sets (with the 9C and 13O beams) and found to be consistent. The results shown by the blue data points represents the average of the two data sets with the error bars indicating their spread.
The correlations for the three 2 emitters are strikingly similar suggesting that the new state is also a prompt 2 emitter. All three distributions, Fig. 4(c), are peaked close to 0.5 where the two protons are emitted with the same energy. This is a characteristic feature of prompt 2 decay as it maximizes the product of their barrier penetration factors. This feature is observed in the correlations of all other ground-state 2 emitters [Egorova:2012, Webb:2019a, Brown:2014b]. Having a maximum at =0.5, also implies that the two protons experience the same barrier and therefore are emitted with the same value. This indicates that the 8.4-MeV state has positive parity consistent with the knockout of a -wave proton in the 9C data set.
Substantial sequential decay of the newly found state, through the known 5/2 intermediate resonance in 7Be, can be more definitively excluded from the invariant-mass distribution of the two -6Li sub events in each 2-6Li event. The data points in Fig. 4(a) show this distribution. The magenta curve shows the expectation for sequential 2 decay obtained using the -matrix formalism [Lane:1958]. This distribution exhibits a peak at 1.5 MeV (from the second sequential step) which is not observed in the data. This peak would also exist in the distribution shown in Fig. 4(c), but as the 8B peak is wide, dividing by smears it out.
For the data set with the 9C beam, an invariant-mass peak for +3He+ exit channel was found with =8.23(18) MeV and =1.30(33) [CharityLong:2025] which are close to the values for the 8.4-MeV peak. This could possibly represents a second decay branch of this 2 emitter. If so, the +3He+ branch has a yield 1.2(4) times that of the 2 branch [CharityLong:2025]. Correlations indicate that the +3He+ exit channel proceeds by an initial 3He+5Li decay and possibly a smaller contribution from +7Be(5/2) decay [CharityLong:2025].
Theory. — We have employed the symmetry-adapted no-core shell model (SA-NCSM) [DytrychLDRWRBB20, LauneyDD16] with the NNLOopt chiral potential [Ekstrom:2013] to calculate the excitation spectrum of 8B and the intrinsic structure of the mass – states of interest. SA-NCSM is an ab initio approach that has been used to successfully predict energy spectra [Sargsyan_A8, heller2022new], electric quadrupole moments [LauneyMD_ARNPS21] and transition strengths [Henderson:2017dqc, Ruotsalainen19, PhysRevC.100.014322], beta decays, clustering features [Sargsyan_A8, DreyfussLESBDD20, DreyfussLTDBDB16], and reaction dynamics [Mercenne:2022, LauneyMD_ARNPS21, BurrowsSumRules], for nuclei up to the calcium region. SA-NCSM yields identical results to the traditional no-core shell model [NavratilVB00, BarrettNV13] for a given inter-nucleon interaction.
Discussion. — Based on the SA-NCSM calculations, candidate levels that can be produced via -wave proton knockout from the 9C beam (0+, 1+, 2+, and 3+) and located in excitation close to the newly observed 2 emitter are the 1, 2 and the 3 states [CharityLong:2025]. These calculations largely reproduce the known levels at lower excitation. The single-proton decay energies of these states to the ground (3/2-) and first excited (1/2-) states in 7Be are more than twice the height of the Coulomb plus centrifugal barrier. In order for a 2 decay branch to complete, the 1 channels must be associated with very small spectroscopic factors. The SA-NCSM spectroscopic factors for the candidate states to these low-lying states in 7Be (Table 1) are indeed quite small (.
The model calculations indicate that the three candidates have significant spectroscopic factors for proton decay to either of the two higher-lying adjacent 5/2- states in 7Be [CharityLong:2025]. Experimentally, the 5/2 state decays almost exclusively to the 3He+ channel [Tombrello:1963, Spiger:1967] while the 5/2 state decays predominantly to the +6Li(g.s.) channel [McCray:1963] which is consistent with the SA-NCSM calculations [CharityLong:2025] and recent Gamow shell-model results [Fernandez:2023].
The 2 and 3 states have large spectroscopic factors for decay to the 5/2 level of 7Be (1.08 and 0.81, respectively) and thus the decay of these levels should populate the +3He+ final state and possibly could help explain the invariant-mass peak observed for that channel at the same value of E∗. The 1 level is predicted to have a significant spectroscopic strength for decay to the 5/2 proton resonance in 7Be and thus could populate the 2+6Li final exit channel via a sequential 2 decay. However, SA-NCSM calculations show that the proton width through the 5/2, while not negligible, is significantly smaller than the width. This difference arises from the fact that prompt decay can emit two -wave isovector paired protons (practically inaccessible to and ), whereas a sequential 2 decay involves a -wave proton emission. This corroborates the experimental results, namely, such a sequential decay is not seen in the experimental correlations [Fig. 4(a)]. This suggests that the decay of 1 to 6Li is similar to the ground-state 2 emitters 6Be [Egorova:2012] and 8C [Charity:2011] where the -wave proton pair channel is accessible and in competition with -wave sequential decay.
When we consider the dominant deformation configurations available from the SA-NCSM calculations, it becomes evident that the 1 state does in fact have a high degree of similarly to the other two =8 2 emitters [8B(0) and 8C(0)] and hence is the best candidate for the new 2 decay (Fig. 5a). The wavefunctions of the parent 2 emitters are dominated by an oblate configuration while those of their daughter states [6Li(1), 6Li(0), and 6Be(0)] are dominated by a prolate configuration. Thus the protons are emitted from the equatorial regions.
In addition, to have a 2 decay unaffected by potential intermediates, a significant overlap between the total orbital angular momentum () of the initial and final states should exist [Fig. 5(b)]. With the 6Li ground state being predominantly =0, and if we assume both protons are emitted from or orbitals, only 8B states with =0, 1, or 2 components should be considered as possible candidates. The 3 assignment can be ruled out as it is predominantly =3 (79%). The 2 state has a significant =2 projection (39%) so it a plausible assignment. The best candidate remains the 1 level which is predominantly =0 (87%), with a moderate =2 projection (11%).
The 2 decay of the =2 members of the triplet are transitions and thus the two emitted protons have zero total angular momentum. With the 1 assignment, the new case is a transition allowing for similar angular-momentum coupling of the two protons thus leading to similar correlations. This is not true for the other candidate spin assignments where the spin couplings of the protons would be completely different. However, the correlations in the triplet are not expected to be completely identical as their decay energies differ: =1.3 MeV (0 8B), 2.1 MeV (0 8C), and 2.7 MeV (1 8B). Subtle differences could also result from the different relative locations and widths of the intermediate states.
The possibility that the 1 state has a significant decay branch to the +3He+ exit channel is not supported by the SA-NCSM calculations for which the 3He+5Li and +7Be(5/2) decay branches are minor [CharityLong:2025].
The same structure should be present in the levels of the mirror nuclei. The mirror 1 level in 8Li (Fig. 1) has been studied via -delay triton emission of 8He. Borge et al. [Borge:1993] measured the energy spectrum, which was fit with a +5He++ sequential decay model, giving an excitation energy of 9.3(1) MeV. This decay path is the mirror of the 3He+5Li path which dominates the state observed in the +3He+ channel observed at almost the same energy. This adds support to the contention that we observed the 3He+5Li decay branch of 8B(1), however, there are other interpretations of the -energy spectrum. Barker’s analysis, allowing for the possibility of more levels and channels, deduced E∗=9.02-9.92 MeV [Barker:1996]. Grigorenko et al. using a three-body ++ model showed that the triton spectrum could be fit without introducing correlations, i.e. with a simple phase-space approach, and found E∗=8.5-8.8 MeV [Grigorenko:1996]. Clearly the deduced excitation energy, of the 1 level in the mirror 8Li, is model dependent. The SA-NCSM and the Greens function Monte Carlo calculations [Wiringa:2000] favor excitation energies at the lower end of the range of these three analyses [CharityLong:2025].
All three analyses of the triton spectrum agree that the (GT) value for the 1 state is very large, in fact one of the largest known. This indicates that the giant Gamow Teller resonance is concentrated in this 1 state and that this state has a structure similar to 8He(g.s.) (cf. Ref. [DytrychLDRWRBB20] for the 8He deformation). Grigorenko et al. refer to it as a “Halo Analog State” [Grigorenko:1996]. This title is also appropriate for our triplet of 2 emitters and for their daughter states which are analogs of 6He(g.s.). While the overall similarity among the parent (daughter) halo analog states is striking, their detailed structure and asymptotics differ, largely due to the Coulomb force.
Conclusions. — The nature of the =8.4-MeV resonance state in 8B observed in [CharityLong:2025] that decays to the 2+6Li channel has been established and shown to have an unexpected resemblance to neighboring resonances. Momentum correlations between the decay fragments indicate that this is a prompt 2 emitting state. Based on comparison to ab initio symmetry-adapted no-core shell-model calculations, this level is assigned as the forth 1+ state which has a spatial structure very similar to 8C and its isobaric analog state in 8B (=2 analog states). Indeed, this newly found state and the two just mentioned constitute a triplet of kindred 2 emitters with similar correlations between the decay products and which, via these decays, remove one unit of isospin, leave the spin and parity unchanged, and converts an oblate parent into a prolate daughter.
Acknowledgements.
This work is supported by the U.S. Department of Energy, Office of Science, Office of Nuclear Physics under Awards No. DE-FG02-87ER-40316, DE-SC0023532, and under the FRIB Theory Alliance award DE-SC0013617. This work is also supported in part by the National Nuclear Security Administration through the Center for Excellence in Nuclear Training and University Based Research (CENTAUR) under grant number DE-NA-0004150. This work benefited from high performance computational resources provided by LSU (www.hpc.lsu.edu), the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility at Lawrence Berkeley National Laboratory operated under Contract No. DE-AC02-05CH11231, as well as the Frontera computing project at the Texas Advanced Computing Center, made possible by National Science Foundation award OAC-1818253.End Matter
Spectroscopic factors and decay widths for states of interest, calculated with the SA-NCSM with continuum [LauneyMD_ARNPS21], are listed in Table 1.
| Channel | 1 | 2 | 3 | |
| SF | ||||
| +7Be(1/2-) | 0.03 | 0.06 | 0.01 | |
| +7Be(3/2-) | 0.02 | 0.02 | 0.08 | |
| +7Be(5/2) | 0.01 | 1.08 | 0.81 | |
| +7Be(5/2) | 0.69 | 0.08 | 0.11 | |
| +7Be(7/2-) | 0.001 | 0.001 | 0.25 | |
| +7Be(5/2) | 0.8(4) | 0.6(3) | ||
| +7Be(5/2) | 0.3(2) | |||
| +6Li(1) | 0.8(2) | 0.04(2) | 0.02(1) | |